49. Thompson Co. has 75,000 shares of $4 par value stock outstanding (total legal capital of...

49.

Thompson Co. has 75,000 shares of $4 par value stock outstanding (total legal capital of $300,000). If Thompson splits the stock 2-for-1, what will be the total legal capital of all Thompson Co. shares after the split?

a. $150,000

b. $450,000

c. $300,000

d. $600,000

50.

A debit balance in Allowance for Uncollectible Accounts indicates that

a. the actual amount of uncollectible accounts was less than the company estimated they would be

b. the company uses the percent of credit sales method to allow for uncollectible accounts

c. the actual amount of uncollectible accounts was more than the company estimated they would be

51.

At Johnson Co., the cost of goods available for sale during the period was $325,000. The total retail selling price for those goods was $500,000. The company had retail sales of $400,000 during the period. What is the estimated cost of the unsold (ending) inventory, using the retail estimation method?

a. $75,000

b. $185,000

c. $65,000

d. $175,000

52.

Assume that the total issue price for a $500,000 face-value bond issue with semi-annual interest payments has been correctly calculated to be $578,450. The market interest rate for bond issues of similar risk is 6%, compounded semi-annually. How much interest expense will accrue and be recorded at the end of the first 6 month period?

a. $30,000.00

b. $34,707.00

c. $17,353.50

d. $15,000.00
Answer #1

Question 49:

The Legal Capital will not be changed even after the Stock split. So, legal capital after stock split be the same

Option C. $ 300,000

Question 50:

Debit balance in the Allowance for Uncollectible Accounts indicates that the actual amount of allownace be more than the Estimated.

Option C. the actual amount of uncollectible accounts was more than the company estimated they would be

Question 51:

For 500,000 retail sale price - the cost be $ 325,000

For 100,000 retails price the cost will be = 100,000 * 325,000/500,000 = $ 65,000

Option C. $ 65,000 is the cost of ending inventory

Question 52:

Interest will be calculated & paid on the Face value of the bonds but not the issue price.

So, Interest Accrued & recorded for the first six months period be 500000*6% * 6/12 = $ 15,000

Option D. $ 15,000 is the answer